### v2

# Quantifying interdisciplinary synergy in higher STEM education

Abstract
We propose a framework to quantify and utilize interdisciplinarity in science and engineering curricula in the case of university-level higher education. We analyze interdisciplinary relations by standardizing large-scale official educational data in Korea using a cutting-edge large language model and constructing knowledge maps for disciplines of scientific education. We design and evaluate single-field and integrated dual-field curricula by adapting pedagogical theory and utilizing information theory-based metrics. We develop standard curricula for individual disciplines and integrated curricula combining two fields, with their interdisciplinarity quantified by the curriculum synergy score. The results indicate higher interdisciplinarity for combinations within or across closely related fields, especially in engineering fields. Based on the analysis, engineering fields constitute the core structure of our design for curriculum interdisciplinarity, while basic natural science fields are located at peripheral stems to provide fundamental concepts.
Interdisciplinarity,
Curriculum Design,
Higher Education,
STEM,
keywords:
[…]
Research
{abstractbox}
1 Introduction
The increasing complexity of challenges across the social, academic, and industrial domains has focused on interdisciplinary approaches to solving advanced problems [1]. In response to this social demand, higher education institutions, e.g., universities, are progressively introducing and expanding their multidisciplinary programs [2, 3, 4, 5]. Consequently, researchers are actively working on the interdisciplinary theory and its curriculum models nowadays [2, 6]. Unfortunately, however, most of such research relies on qualitative assessments based on theoretical arguments or small-scale data, which makes its versatile application questionable.
The primary objective of this research is to establish a principled data-driven framework for designing and evaluating interdisciplinary-integrated curricula. To achieve this, we analyze large-scale real-world data extending beyond traditional curriculum design approaches. We develop a principled data processing pipeline and a quantitative framework to measure interdisciplinarity across large-scale academic curricula. Our methodology begins with standardizing extensive scientific and technological natural language data. We build a standardized dataset using cutting-edge natural language processing (NLP) techniques, such as large-scale language models (LLMs) [7], and higher-order methods [8]. We link this dataset to departments across institutions through a bipartite network to construct knowledge maps. By projecting this network, we analyze standard subjects’ relationships and interconnections among various disciplines.
In particular, we focus on the science, technology, engineering, and math (STEM) disciplines. The decision to concentrate on the STEM disciplines in our analysis is largely influenced by the dynamic evolution and heightened demand these disciplines are experiencing, driven by recent technological advancements. The STEM disciplines are at the forefront of innovation, playing a crucial role in addressing some of the world’s most pressing challenges. As a result, there is an increasing demand for professionals skilled in these areas, necessitating a stronger emphasis on interdisciplinarity to equip future experts with the ability to integrate knowledge and skills from various domains [9]. Recent policy directions, including those led by the U.S. National Science Foundation, emphasize the necessity of interdisciplinary training in STEM to prepare the next generation of leaders for complex, real-world challenges [10]. New policies are advocating for an interdisciplinary approach within STEM education, recognizing that the complexity of modern challenges requires solutions that are not confined to a single discipline [11]. By focusing on STEM, this research aims to contribute to the development of educational frameworks that not only respond to these policy shifts but also prepare students to thrive in a rapidly evolving technological landscape.
Leveraging information theory [12] to show the difference between the major distributions quantitatively, we propose a single-discipline standard curriculum model that reflects real-world constraints such as time, resources, and institutional policies. In addition, we hypothesize integration majors, which combine two distinct fields, and develop a double-field integrated curriculum for each possible combination. To assess the interdisciplinarity of these integration majors, we present a measurement framework taking both the curriculum similarity and the synergy into account. This metric allows us to determine whether a given double-field integrated curriculum achieves true interdisciplinarity or merely a multidisciplinary combination by systematically evaluating all possible combinations and comprehensively analyzing their degree of integration.
The paper is organized as follows. In Sec. 2, we introduce our data and the process of course standardization with the large language model (LLM). Based on the processed data, we first analyze the structural properties of the department similarity in terms of curricula in Sec. 3. As the main topic of our work, we present the interdisciplinary curriculum model and the quantitative measure of synergy between disciplines in Sec. 4. We summarize our work and discuss its implications in Sec. 5.
2 Data processing of course standardization
For this research, we develop a data processing framework that allows us to assess the quality of intermediate steps and final results across sectors and scales. To present higher education boundary’s standard single-field curricula and double-field integrated curricula and quantify their multidisciplinarity and interdisciplinarity, we employ datasets primarily sourced from the “Higher Education in Korea” portal publicly operated by the Ministry of Education in Korea for the 2024 spring semester. The data sets are officially collected, processed, and validated by the Ministry of Education in Korea as a national census of higher education, so it is highly authoritative and trustworthy. Our primary data set includes the “Curriculum per Educational Unit by School” for the 2024 spring semester [13]. In addition, we utilize the “List of Undergraduate Departments and Majors by School” for the same semester [14], as metadata to identify departments and educational fields. These data sets were integrated to extract STEM departments’ curricula from departments classified under the “Standard Classification Information (표준분류정보 in Korean)” system [15].
As introduced in Sec. 1, we focus on the departments running four-year undergraduate programs within STEM fields classified under Engineering and Natural Sciences in the Standard Classification (the highest-level classification). Specifically, within Natural Sciences, we extracted data from departments categorized as “Basic Natural Sciences,” which include “Chemistry, Life Science, and Environment” and “Mathematics, Physics, Astronomy, and Earth Science” in the Standard Classification (the intermediate-level classification). As a result of our preprocessing and comprehensive data integration, we constructed a standardized dataset comprising curricula from 2 84128412\,8412 841 STEM departments [^1] The original data included 2 85128512\,8512 851 unique STEM undergraduate departments. However, due to military security restrictions and data structure errors, only 2 84128412\,8412 841 departments were ultimately analyzed. across $161$ higher education institutions, including a total of 126 437126437126\,437126 437 raw course entries. We use departments as the primary unit of analysis because our data, sourced from the Ministry of Education in Korea, is organized at this level, providing a consistent framework with around $80$ credits (for major courses) over four years per department. In Korea, departments serve as the de facto units of disciplines due to the structured nature of higher education, aligning with current educational practices and government initiatives that encourage interdisciplinary integration. While finer distinctions at the program level might be desirable, current data limitations necessitate this approach. Future research could explore more granular analyses as detailed data becomes available, such as through “microdegree” programs aimed at fostering small units of interdisciplinary education.
We also note that different departments may offer courses with varying credit allocations, but the overall credit requirement standardizes the instructional volume, facilitating consistent cross-departmental comparisons. All curricula are assessed under a rather uniform constraint in the total number of credits, ensuring that variations in course numbers do not bias our analysis. This approach is again consistent with Korea’s centralized higher education policy, characterized by a national standard for four-year undergraduate curricula. Although such discrepancies might have greater implications in countries with more flexible curricula, their impact would be minimal in the Korean educational context.
The courses are obviously building blocks of the coursework in each department. However, we cannot simply use the raw course data as it is, because there are multiple nomenclatural variants of essentially the same course, e.g., “quantum mechanics” and “quantum physics” in physics-related departments. Therefore, it is crucial to standardize different variants of course names in the raw data to accurately identify course units that are as meaningfully independent as possible shared by different departments and institutions. Without the process, only the courses with exactly the same name would be identified as the same entity. Such a task in this scale would be next to impossible manually, but cutting-edge LLMs enable us to do this task systematically. In particular, we use Gemini-1.5 Pro [7], which is commonly referred to as one of the most accurate language models as of July 2024, when we performed this part of data processing. By employing higher-order data processing techniques combined with the LLM, which will be discussed in Sec. 2.2, we systematically process these raw courses into 24 4762447624\,47624 476 standardized courses. Our approach enables us to effectively process large-scale datasets while accurately capturing the semantic complexities inherent in natural language and disciplinary knowledge.
Figure 1: A schematic illustration of education data processing with LLM. (a) The tokenizing process of course data by LLM (Sec. 2.1). Each course is broken into distinct tokens, marked by different colors. (b) For each token, the courses including the token are listed below. Among them, the courses identified as the same by the LLM are connected in the form of hyperedges (enclosed by the same connected shaded area). (c) Each connected component constructed by collecting the hyperedges in the process described in the panel (b) is identified as each standardized course.
2.1 Tokenizing from the raw course data with LLM
Even with the LLM, extracting the standardized courses directly from the exhaustive list of 126 437126437126\,437126 437 raw courses is not plausible due to the combinatorial explosion. Our strategy to circumvent this problem to proceed is the following. We first utilize tokenization to segment each of the raw course names into meaningful components that serve as the basic indicators for further processing. Each course $C_{i}$ is decomposed into a set of tokens $\{t_{1},t_{2},\dots,t_{m}\}$ as illustrated in Fig. 1(a). To ensure accurate processing of the natural language in the raw course names (Korean), we use prompts tailored to its linguistic properties with specific examples (prompt 1 in Appendix A). Through this tokenization process, we obtain 8 84188418\,8418 841 unique tokens from the raw course data. By focusing on the shared tokens among different courses, this preprocesses the data to highlight potential semantic similarities for further analysis. To process raw courses effectively, for each token, we collect the list of courses including it as illustrated in Fig. 1(b).
2.2 Constructing hyperedges among raw courses with LLM
Now we have the set of raw courses for each token $t$ as in Fig. 1(b), and our next step is to collect the effectively same raw courses as intermediate-level units. Mathematically, this step involves identifying higher-order relationships (i.e., hyperedges) [8] among raw courses—beyond simple pairwise associations—so as to capture more nuanced semantic groupings that reflect disciplinary structures [16, 17, 18]. Similar to the tokenization process described in Sec. 2.1, we again use a prompt that effectively clusters identical courses based on shared topics or subfields (prompt 2 in Appendix A), depicted in Fig. 1(b).
While conventional NLP techniques have long been applied to large-scale textual data analysis, they often rely on rigid pipelines tailored to narrow tasks. In contrast, we adopt large language models (LLMs) due to their superior contextual understanding and adaptability. Recent studies have demonstrated that LLMs consistently outperform traditional NLP tools in extracting nuanced semantic relationships from domain-specific scientific corpora [19]. Their ability to integrate context, handle ambiguous terminology, and operate with minimal labeled data makes them especially effective for tasks involving the clustering of STEM course descriptions, which are rich in discipline-specific language and complex structural patterns.
Through this LLM-based higher-order clustering process, we generate 91 9649196491\,96491 964 hyperedges from all tokens, which encapsulate the intermediate-level interrelationships present in the raw course data.
2.3 Standardizing courses
Our final step to standardize courses is based on the list of hyperedges in Sec. 2.2. As illustrated in Fig. 1(c), we merge all of the courses inside each of the connected components. Intuitively, this definition captures not only directly connected courses within a hyperedge but also courses that are indirectly linked through the chain of associations. For example, if course $C_{1}$ is associated with hyperedge $H_{k}$ and course $C_{2}$ is linked to both $H_{k}$ and $H_{l}$, then $C_{1}$ and $C_{2}$ belong to the same connected component. The set of connected components constitutes our final set of standardized courses $\mathcal{S}$. In this way, each connected component groups the raw courses that are semantically or structurally similar based on shared tokens or associations identified via the hyperedge construction process.
Through this standardization process, we reduce 126 347126347126\,347126 347 raw courses to 24 4762447624\,47624 476 standardized courses from STEM departments of Korean universities. This reduction merges redundant or overlapping courses into unified representations while preserving their semantic integrity and higher-order relationships, for more interpretable and robust analysis. By our processing, the standardized courses provide a deeper understanding of the underlying structures and interconnections within curricula [20]. For a concrete example, we provide the step-by-step process in the case of statistical physics in Appendix B.
3 Department similarity based on standardized courses
Figure 2: A schematic illustration of the DSN constructing process based on standardized course data. The vectors on the left represent the credits for each standardized course for multiple departments across institutions, where each element corresponds to the total weight (the credit sum) assigned to a specific standardized course. Each row represents a department, and each column corresponds to a standardized subject. The right matrix represents the pairwise similarity $S_{ij}$ according to Eq. (1). The resulting similarity matrix is visualized as a heatmap, where darker shades indicate higher similarity between departments.
Figure 3: Alluvial diagram showing the membership reorganization of community structures at four representative resolution values: $\gamma=0$, $0.56$, $1.12$, and $1.77$, where the community structures show global consistency (see Appendix C for details), where the vertical positions of the nodes are determined to minimize crossing. At $\gamma=0.56$, each node and link represents the community’s aggregated departmental leanings (PSE: red, TSE: green, and MLSE: blue), and at finer resolution values, the weighted sum of colors represents the reorganization of departmental classifications.
To verify the validity and utility of the standardized courses extracted in Sec. 2, we constructed a department similarity network (DSN), where nodes represent departments, and the weighted edges represent their department similarity in terms of their curricula (Fig. 2[21, 22]. The department similarity is quantified using the weighted Jaccard similarity coefficient, defined as:
$S_{ij}=\frac{\displaystyle\sum_{k=1}^{n}\min(c_{ik},c_{jk})}{\displaystyle\sum_{k=1}^{n}\max(c_{ik},c_{jk})},$ (1)
where $c_{ik}$ represents the credit-weighted value of the standardized course $k$ (the credit of a standardized course is defined as the sum of the credits assigned to the raw courses corresponding to the standardized course) in department $i$, and $S_{ij}$ measures the proportion of standardized coursework shared between departments $i$ and $j$.
We explore the structural properties of DSN, in the viewpoint of communities [23, 24] of various scales. We take the popular modularity-maximization algorithm Louvain [25] with the tunable resolution parameter $\gamma$. We choose four representative values of $\gamma=0$, $0.56$, $1.12$, and $1.77$, where the overall community structures show global consistency among the results (Appendix C[26]. We illustrate the scale-dependent reorganization of communities across the four resolution values as an alluvial diagram in Fig. 3. Each block in the diagram corresponds to a community within DSN at a particular $\gamma$ value, with the inter-connection between the blocks in the adjacent $\gamma$ values indicating the flow of department membership.
At $\gamma=0$, the DSN consists of a single community by the mathematical property of the modularity function [23, 24], marked by gray. As the resolution increases to $\gamma=0.56$, the network is divided into three main distinct communities: (1) pattern science and related engineering (PSE) such as physics, mathematics, and software engineering (the top red block), (2) technical systems engineering (TSE) such as architectural engineering (the middle green block), and (3) molecular, life science, and related engineering (MLSE) such as chemistry, biology, and chemical engineering (the bottom blue block). Departments that do not belong to these main clusters correspond to the thin gray block. It is interesting to note that the first meaningful separation of communities is not the division between natural science and engineering that one might imagine, but that between PSE and MLSE, where different types of sciences mixed with various engineering departments for both, in addition to TSE.
As we increase the resolution parameter further, at an intermediate resolution value $\gamma=1.12$ and the finest resolution $\gamma=1.77$, the aforementioned three communities are refined further. The colors of the communities at these resolutions are adjusted to reflect their original affiliations at $\gamma=0.56$. When departments from different communities are remerged into a single community at a finer resolution value, their colors are blended to highlight the interdisciplinary composition. Note that as the resolution increases, the red (PSE) and green (TSE) communities are fragmented into smaller communities. In contrast, the blue (MLSE) communities retain a significant portion of their original nodes, reflected by similar block sizes even at finer resolution values.
One can track the membership reorganization of specific departments across different scales, and it would be interesting to see the flow of less traditional and interdisciplinary departments. For example, in the case of the Department of Scientific Computing at Pukyong National University, whose curriculum is highly interdisciplinarity with the studies on scientific computing from various disciplines, belongs to a computer-science-related community at intermediate resolution values. However, as the resolution gets finer, this department transitions into the physics-related community presumably because of its interdisciplinary curriculum.
4 Interdisciplinary curriculum design and quantifying its synergy
Figure 4: A schematic diagram of our curriculum design process. (a) A recap of the weighted bipartite network between departments and standardized courses, described as a weighted matrix in Fig. 2. The detailed course standardization process is stated in Sec. 2. (b) The standardized-course-mode projection from the bipartite network, from which the curriculum is aggregated by sequential addition of candidate courses based on the information-theoretical criterion described in panel (c).
In this section, based on the standardized courses, we suggest a framework for designing curricula for potential interdisciplinary fields, followed by evaluating the synergy from such field combinations. First, we need to set up a list of standard academic disciplines. As a reminder, we have constructed a comprehensive list of STEM curricula in terms of standardized courses for each department in Sec. 2 and Sec. 3. This process involves constructing a bipartite network with nodes representing 24 4672446724\,46724 467 standardized courses and 2 84128412\,8412 841 STEM departments. The edges in the bipartite network are weighted by the total credit hours associated with each course [Figs. 2 and 4(a)]. To quantify the relationship between the standardized courses, we project the bipartite network into the unipartite network composed of standardized courses, whose edges are weighted by the sum of credits in the departments sharing the course pairs [Fig. 4(b)].
4.1 Single-discipline curriculum model
From the original source data provided by Ministry of Education in Korea [15], the classification of “Standard Classification information (the lowest-level classification)” is designated as the list of $51$ independent disciplines [^2] Among the $52$ disciplines in the data set [15], we exclude the “not classified elsewhere (N.C.E.)” entry for proper analysis. . For each discipline, we define the standardized credits of the standardized courses in the discipline as the most frequent nonzero credits in the departments belonging to the discipline. For instance, if the credits of a standardized course $\mathcal{C}$ appear as $\{6,6,3,6,4,0,0,0,0,0\}$ in $10$ departments corresponding to a discipline $\mathcal{D}$, the standardized credit of $\mathcal{C}$ in discipline $\mathcal{D}$ is $6$. From now on, “courses” and “credits” always refer to standardized ones, unless stated otherwise.
First, we need to design a curriculum model for a single discipline before moving on to the interdisciplinary curriculum. To design a standard single-field major curriculum for each discipline, we begin with five “seed” courses as the foundation, by choosing the five courses with the maximum credit assigned to the discipline [^3] In the case of ties, we choose a desired number of courses uniformly at random—the missing courses by chance are likely to be included at the initial stage of model, so it does not make meaningful difference.. The “aggregation” process of course attachment is composed of the two parts: sorting the candidate courses in the descending order of the weight sum between each course with the currently existing courses in the course-projected network [Fig. 4(b)], deciding whether we accept the candidate course (from the interactively updated queue of the sorted list) in the curriculum or not, and repeating this trial until we reach a certain threshold. A state of a curriculum during the aggregation process is described by the $n$($=|\mathcal{S}|$)-dimensional vector $\mathbf{w}_{\mathrm{model}}=\left(w_{1},w_{2},\cdots,w_{n}\right)$, where each index corresponds to a course and its corresponding element $w_{i}$ represents the course $i$’s credit in the discipline.
To evaluate the degree of new information introduced by the interdisciplinary curriculum, we compute the Jensen-Shannon divergence (JSD) between the course-weighted vectors of the simple union and the integrated model. We denote this divergence as $\mathcal{J}$, which provides a symmetric and bounded measure of dissimilarity between the two distributions:
$\mathcal{J}\left(\mathbf{w}_{\mathrm{model}},\mathbf{w}_{d}\right)=\frac{1}{2}\left[D_{\text{KL}}\left(\mathbf{w}_{\mathrm{model}}\middle\|\mathbf{m}_{d}\right)+D_{\text{KL}}\left(\mathbf{w}_{d}\middle\|\mathbf{m}_{d}\right)\right]\,,$ (2)
where $\mathbf{m}_{d}$ is the element-wise average vector of the two vectors:
$\mathbf{m}_{d}=\frac{1}{2}\left(\mathbf{w}_{\mathrm{model}}+\mathbf{w}_{d}\right)\,,$ (3)
and the Kullback-Leibler (KL) divergence $D_{\text{KL}}$ measures the directed discrepancy between a given weight vector $\mathbf{w}$ and a reference vector $\mathbf{m}$,
$D_{\text{KL}}\left(\mathbf{w}\middle\|\mathbf{m}\right)=\sum_{i=1}^{n}w_{i}\log\left(\frac{w_{i}}{m_{i}}\right)\,.$ (4)
The process sequentially adds candidate courses similar to existing department curricula in the discipline, on average. At each step, the single-field average JSD is calculated as:
$\langle\mathcal{J}\rangle_{\mathcal{D}}=\frac{1}{|\mathcal{D}|}\sum_{d\in\mathcal{D}}\mathcal{J}\left(\mathbf{w}_{\mathrm{model}},\mathbf{w}_{d}\right)\,.$ (5)
The addition of the candidate course in the model is accepted if $\langle\mathcal{J}\rangle_{\mathcal{D}}\leq 0$, and rejected if $\langle\mathcal{J}\rangle_{\mathcal{D}}>0$. Finally, we terminate the model aggregation when the sum of the credits reaches at least $80$, which is a typical number of credits for major courses in four-year undergraduate degrees in Korea. The optimized curriculum of discipline $\mathcal{D}$, denoted by $\Omega\left(\mathcal{D}\right)$, can be expressed as the credit-weighted vector for each course
$\tilde{\mathbf{w}}_{\mathcal{D}}=\left(\tilde{w}_{\mathcal{D};1},\tilde{w}_{\mathcal{D};2},\cdots,\tilde{w}_{\mathcal{D};n}\right)\,,$ (6)
where $\tilde{w}_{\mathcal{D};i}$ is the credit of course $i$ in the optimized curriculum of discipline $\mathcal{D}$. The vector for each discipline is standardized by the optimization process described in Sec. 2, resulting in the uniform credit sum $\approx 80$, to ensure comparability and consistency in the curriculum design process.
4.2 Interdisciplinary curriculum model
To construct a double-field interdisciplinary curriculum model, we start with a seed of five courses in each field and their assigned credits for the $10$ courses in total. Similar to the single-discipline model in Sec. 4.1, the candidate courses for addition are interactively sorted by the sum of weights [in the projected network in Fig. 4(b)] attached to the current courses in the model.
In this case, the JSD for the acceptance criterion becomes
$\langle\mathcal{J}\rangle_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}=\frac{1}{|\mathcal{D}_{A}|}\sum_{d_{A}\in\mathcal{D}_{A}}\mathcal{J}(\mathbf{w}_{\mathrm{model}},\mathbf{w}_{d_{A}})+\frac{1}{|\mathcal{D}_{B}|}\sum_{d_{B}\in\mathcal{D}_{B}}\mathcal{J}(\mathbf{w}_{\mathrm{model}},\mathbf{w}_{d_{B}})\,,$ (7)
where $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ are the two respective disciplines.
This iterative process retains only the courses that reduce the average JSD across all of the departments belonging to both disciplines, and the same threshold of $80$ credits applies for the model termination. Similar to the single-discipline case, the optimized interdisciplinary curriculum of disciplines $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, denoted by $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$, can be expressed as the credit-weighted vector for each course
$\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}=\left(\tilde{w}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B};1},\tilde{w}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B};2},\cdots,\tilde{w}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B};n}\right)\,,$ (8)
where $\tilde{w}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B};i}$ is the credit of course $i$ in the optimized interdisciplinary curriculum of disciplines $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$.
4.3 Quantifying interdisciplinarity with the curriculum synergy score
In Sec. 4.2, we have proposed an information-theoretical model to evaluate the degree of integration achieved by interdisciplinary curricula. The central aim is to design a curriculum that reflects both semantic cohesion across departments and pedagogical completeness within a combined field of study. To formalize this, we define an optimized curriculum set $\Omega(\mathcal{D})$ for any given set of departments $\mathcal{D}$. The selection process is based on a semantic filtering criterion: each course must contribute to reducing the average divergence in Eq. (5) across departmental curricula. The credit constraint mentioned in Sec. 4.1 ensures that the resulting curriculum maintains the completeness required for a full major according to national education standards. An analogous formulation is applied to interdisciplinary combinations $\mathcal{D}_{A}\otimes\mathcal{D}_{B}$. Based on this formal construction, we introduce the curriculum synergy score (CSS), which evaluates whether the algorithmically constructed interdisciplinary curriculum $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$ achieves a deeper level of integration than a simple combination of the two disciplines, denoted by $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$.
Although the concept of synergy is nontrivial, our operationalization focuses on two key aspects: (1) the interdisciplinary curriculum should maintain a sufficient degree of alignment with its constituent disciplines; and (2) it should generate new semantic content that is not trivially inherited from the union of its parts. This second aspect is grounded in the notion of “surprisal” from information theory [12], reflecting the emergence of integrative knowledge structures not found in either discipline alone. We assume the following course-weighted vector for the double-major,
$\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}}=\left(\max\left\{\tilde{w}_{\mathcal{D}_{A};1},\tilde{w}_{\mathcal{D}_{B};1}\right\},\cdots,\max\left\{\tilde{w}_{\mathcal{D}_{A};n},\tilde{w}_{\mathcal{D}_{B};n}\right\}\right)\,,$ (9)
which is simply the element-wise weighted union of the optimized curricula in Eq. (6) for the two disciplines $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$.
Let us explore the mathematical interpretation of integrated disciplines through the lens of pedagogical theory. In contrast to multidisciplinarity, where disciplines coexist independently without significant integration, interdisciplinarity focuses on the meaningful synthesis of knowledge, tools, and perspectives from multiple fields [27, 28]. Achieving interdisciplinary synergy requires a substantial overlap between disciplines, as opposed to merely combining them in a multidisciplinary manner. Our JSD-based CSS is specifically designed to capture this distinction by quantifying how effectively an integrated curriculum encompasses both shared similarities and the emergence of new information, as compared to a basic multidisciplinary curriculum. This approach helps distinguish truly integrative designs from mere aggregations by evaluating whether the synthesized curriculum provides added pedagogical value rather than duplicated content. For students, a high CSS implies a more coherent learning experience, minimizing redundancy and promoting cross-disciplinary connections that are more than the sum of their parts. To quantify the synergy reflecting those aspects, we employ an information-theoretic method based on prior research in educational fields [4, 5, 29, 30, 31, 32]. For the new information emerging from the interdisciplinary curriculum, we again apply the JSD for the course-weighted vectors to quantify the informational distance between $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$ and $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$, given by
$\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)=\frac{1}{2}\left[D_{\text{KL}}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}}\middle\|\mathbf{m}_{\cup\otimes}\right)+D_{\text{KL}}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\middle\|\mathbf{m}_{\cup\otimes}\right)\right]\,,$ (10)
where $\mathbf{m}_{\cup\otimes}$ is the element-wise average vector of the two vectors:
$\mathbf{m}_{\cup\otimes}=\frac{1}{2}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}}+\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)\,,$ (11)
and the KL divergence $D_{\text{KL}}$ between two vectors is defined previously in Eq. (4). Note that in contrast to the JSD for designing the curriculum, the JSD here measures the amount of new information between $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$ and $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$, so larger values indicate the possibility of unexpected synergy as a result of interdisciplinary curricula. Of course, as we previously mentioned, one cannot simply use the JSD alone because the “unexpected” information can be caused by synergetic or completely unrelated combinations. Therefore, we have to set an extra factor to filter those unrelated combinations.
Our strategy for filtering is to use a similarity measure between the disciplines of interest. We use the following simple measure that quantifies the degree of shared courses between the two disciplines, measuring the extent of overlap in their optimized weighted course vectors. It is computed as
$\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}=\frac{\displaystyle\sum_{i=1}^{n}\min\left\{\tilde{w}_{\mathcal{D}_{A};i},\tilde{w}_{\mathcal{D}_{B};i}\right\}}{\displaystyle\sum_{i=1}^{n}\max\left\{\tilde{w}_{\mathcal{D}_{A};i},\tilde{w}_{\mathcal{D}_{B};i}\right\}}\,.$ (12)
Finally, we define the CSS between disciplines $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ as the product of the two factors in Eq. (10) and Eq. (12),
$\Theta\left(\mathcal{D}_{A},\mathcal{D}_{B}\right)=\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}\,,$ (13)
which takes both the new information from the interdisciplinary curriculum design and the discipline similarity into account. For instance, if two disciplines are too similar, $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}\approx 1$ but $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)\approx 0$ and thus $\Theta\left(\mathcal{D}_{A},\mathcal{D}_{B}\right)\approx 0$. Another extreme case is the two totally unrelated disciplines, which may induce a large value of $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ but $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}\approx 0$ and thus $\Theta\left(\mathcal{D}_{A},\mathcal{D}_{B}\right)\approx 0$ again. An ideal interdisciplinary combination for synergy corresponds to a subtle balance between emergent information and decent similarity.
Figure 5: (a) The histogram of similarity $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ between discipline pairs. (b) The histogram of information distance $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ between the double-major and interdisciplinary curriculum models for discipline pairs. (c) The scatter plot between them.
The CSS, therefore, quantifies the synergy between disciplines that cannot solely be captured by simple similarity measures, which highlights the added value of proper interdisciplinary integration. We observe the distribution of each factor separately, and one can check that the vast majority of $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ values are close to 00 [Fig. 5(a)], which is understandable by the fact that most discipline pairs would be quite different to each other. On the other hand, the distribution of $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ shows the most prominent peak around $0.3$ presumably corresponding to the overall scale difference between $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}}$ and $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}$, and the second tallest peak around $0.6$ [Fig. 5(b)]. There is no notable correlation between $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ and $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ [Fig. 5(c)], which also justifies the formalism of $\Theta\left(\mathcal{D}_{A},\mathcal{D}_{B}\right)$ as the product of them. In Appendix D, we provide the plots in the double-logarithmic scale in Fig. 9 to help the reader better understand the range and tail behavior of the similarity values.
For notable examples, we provide three examples in supplementary tables: two examples of beneficial synergy (‘physics’ + ‘semiconductor engineering’ in Table S1 and ‘chemistry’ + ‘high polymer engineering’ in Table S2) and an example of no synergy (‘physics’ + ‘chemistry’ in Table S3). It is interesting to note that science disciplines and their related engineering disciplines are much better combinations than two science disciplines, and this tendency will be highlighted again in the analysis presented in the next subsection as well. Other examples include combinations ‘biotechnology + energy engineering,’ which are intuitively synergistic, alongside less conventional yet plausible pairs such as ‘nautical science + civil engineering’ and ‘optical Engineering + semiconductor engineering.’
4.4 The CSS network analysis
The CSS between discipline pairs in Eq. (13) forms a weighted matrix or network representing the level of educational synergy between disciplines. As the number of disciplines is $51$, even a visual inspection compared to the hierarchical structure of disciplines provided by the data set [15] is plausible. We show the $51\times 51$ heatmap of the CSS for all of the STEM discipline pairs in Fig. 6. Overall, large CSS values are prominent among closely related engineering disciplines, hierarchically close to each other according to the standard classification in higher levels (forming dark block diagonal patterns) [15], e.g., the chemical engineering cluster, the computer science cluster, electrical engineering cluster, the mechanical engineering cluster, and the material engineering cluster. In contrast, natural science disciplines do not show large CSS values to each other (the lack of dark block diagonals in the natural science region), and they have large CSS values with discipline-specific engineering counterparts, e.g., (life science $\otimes$ biotechnology) and (physics $\otimes$ semiconductor engineering). This connection pattern is an archetypal mesoscale structural example of core-periphery structures [33], with multiple cores composed of separate modules [34].
Figure 6: The heat map of CSS values between the $51$ disciplines (the lowest-level classification) in the data set. The rows and columns are sorted by the intermediate-level classification shown in the upper hierarchical level drawn on the left [15].
Figure 7: The community (a) and core-periphery (b) structures in the CSS network, where the size of a node (discipline) is proportional to the number of departments in the discipline. (a) The fraction of colors represents the community membership [35] of each node, from $1000$ community ensembles with the resolution parameter $\gamma=1$. (b) Darker colors represent larger core score [33] values for nodes.
To characterize such a mesoscale structure, we try both community [23, 24] and core-periphery [33] detections. For the community structure in this discipline-synergy network, we apply the detection algorithm that allows the overlapping communities (a single node can have multiple community affiliations) [35], which also yields the relative strength of community membership for each node, to emphasize the interdisciplinary relations. The community structures shown in Fig. 7(a) indicate such overlapping communities for disciplines, in particular, drastically contrasting cases of PSE and TSE (belonging to multiple communities) and MLSE (effectively belonging to a single community).
In addition, we verify the aforementioned core-periphery separation by using the core score [33]. As expected from the heat map in Fig. 6, we find that the disciplines classified as basic nature science: physics, chemistry, life science, and statistics, are located in the periphery, as shown in Fig. 7(b). In contrast, the disciplines related to engineering, which are assigned to multiple communities, show high core scores. Therefore, it verifies the existence of multiple cores [34] composed of engineering disciplines synergetic to each other, and the rest of the peripheral parts are composed of natural science disciplines.
One might perceive this result as unintuitive considering the role of basic science and engineering, but in terms of academic curricula, the application side (engineering) is more synergetically attached to other disciplines than basic science, which endows most fundamental concepts and frameworks as basic resources of understanding the entire STEM fields. In other words, the interdisciplinary synergy is not readily obtained due to the unique role of natural science disciplines, which in turn suggests the necessity of maintaining their established curricula as anchoring sources of STEM education, as suggested by the CSS value of zero between physics and chemistry in Table S3. In the case of research collaboration [36], of course, the pattern of interdisciplinarity can be different, so a more comprehensive understanding across education and research can be a natural candidate for future works.
5 Discussion and conclusion
In this paper, based on the official university-level STEM curriculum record provided by the Ministry of Education in Korea, we have proposed a comprehensive process of systematic LLM-based course standardization, the quantitative analysis of the organizational structure of departments based on the curriculum similarity, and the interdisciplinary curriculum design. We have introduced the curriculum synergy score (CSS) to evaluate the performance of the interdisciplinary curriculum model and identify synergetic discipline pairs.
The result from the curriculum-wise organizational structure of departments suggests a deeper division between PSE and MLSE than that between science and engineering, with notable reorganization of less traditional departments depending on the organizational scale. The departments belonging to MLSE are more coherently organized, reflected by their intact community structure at finer resolution, while engineering departments show more flexible relationships within STEM disciplines. Such organizational flexibility of engineering departments is also highlighted in the synergetic relationship quantified by the CSS. The engineering disciplines constitute multiple core structures in the CSS network, while basic science disciplines are located in the peripheral part, with their unique sources indispensable for STEM education.
Considering the result of our data analysis, we call for a more careful approach to designing interdisciplinary education policy. For instance, the fact that basic science disciplines provide fundamental knowledge to understand STEM does not automatically imply that it is beneficial to merge them with other disciplines; rather, we may need to acknowledge their own merit as fundamental sources of knowledge and skills. However, it is important to acknowledge that our algorithmic approach alone cannot overcome complex educational challenges. When our data-driven insights are integrated with evolving educational needs, we can design more effective curricula that better prepare students for future demands. As we have demonstrated that this type of quantitative analysis can bring such insights, A natural extension of this work can be one that includes non-STEM disciplines such as literature, art, history, etc., possibly across different countries. In particular, in this era of technological breakthroughs and innovations in education, the systematic analysis of temporal change in curriculum structures is also highly desirable. On a theoretical side, the interdisciplinary curriculum models involving three or more related disciplines are a candidate for future work. We hope that our work, equipped with nationwide exhaustive data and LLM, provides a nice example for this direction.
Abbreviations
STEM, science, technology, engineering, and mathematics; DSN, department similarity network; CSS, curriculum synergy score.
Declarations
Ethics approval and consent to participate
The data sets utilized in this study have been sourced from Higher Education in Korea (대학알리미 in Korean) [13, 14, 15], a public data service of the Ministry of Education in Korea. This study does not incorporate any personal information.
Consent for publication
All of the authors consent to publication.
Availability of data and material
The original data is available on Higher Education in Korea (대학알리미 in Korean) [13, 14, 15], a public data service of the Ministry of Education in Korea. The processed data, the codes, and the curriculum design are available at https://github.com/Gahyoun/STEM-curriculum-2024-South-Korea-. The full result of our optimization process in the cases of single and interdisciplinary curricula is available in the ‘result’ in Github.
Competing interests
The authors declare that they have no competing interests.
Funding
This research was supported by the National Research Foundation of Korea (NRF) under the grants RS-2021-NR061247 (GG and SHL) and RS-2022-NR069864 (JY).
Author contributions
GG designed the research. GG collected and processed the data. GG analyzed and visualized the data. JY provided LLM and data science technical consultation and advice. GG and SHL designed and refined the CSS. All three authors wrote the paper. All authors have read and approved the final manuscript.
Acknowledgements
The authors thank Jibeom Seo and Beom Jun Kim for their valuable discussions and contributions during the initial stages of this research, and appreciate Daekyung Lee for providing the network-community analysis code.
Appendix A The LLM prompts used in data processing
  1. 1.
    Tokenize the given Korean subject data to reflect the characteristics of STEM data. Example: 열역학 is about 열 and 역학, 통계물리 consists of 통계 and 물리. 전자물리 and 전자물리학 are 전자 and 물리, 화학 is 화학, 전자기학 is 전자기학, 유기생물 is about 유기 and 생물. If the data is in English, translate into Korean and discuss it.
  2. 2.
    For the following list of subjects from the Korean Basic Natural Science Department, translate each subject into English and then assign each to a specific group based on the core topics and content they cover. The categorization ensures high granularity to reflect specific aspects of the subjects. Subjects like 물리학 (physics) and 물리학실험 (physics experiment) are grouped only if they address the same concepts. Closely related subjects such as 고급물리학 (advanced general physics) and 고급현대물리학 (advanced modern physics) are separated if they emphasize different principles.
Appendix B Example of tokenization and course standardization
We provide a step-by-step process of tokenizing and course standardization with an example of four raw course names that might be related to statistical physics. First, the following four raw course names are tokenized as follows, where $T(C_{i})$ refers to the set of tokens for course $i$:
$C_{1}\equiv\text{Statistical Physics:}$: $\quad T(C_{1})=\{\text{Statistics},\text{Physics}\},$
$C_{2}\equiv\text{Statistics:}$: $\quad T(C_{2})=\{\text{Statistics}\},$
$C_{3}\equiv\text{Thermodynamics:}$: $\quad T(C_{3})=\{\text{Thermo},\text{Dynamics}\},$
$C_{4}\equiv\text{Statistical Thermodynamics:}$: $\quad T(C_{4})=\{\text{Statistics},\text{Thermo},\text{Dynamics}\},$
$C_{5}\equiv\text{Statistical Mechanics:}$: $\quad T(C_{5})=\{\text{Statistics},\text{Mechanics}\}.$
For each token $t_{j}$, we can collect the set $\mathcal{G}(t_{j})$ of courses including it, for example,
$\mathcal{G}(\text{Statistics})=\{C_{1},C_{2},C_{4},C_{5}\},\quad\mathcal{G}(\text{Thermo})=\{C_{3},C_{4}\}.$
After tokenization, higher-order clustering identifies semantic relationships among the courses. Using an LLM, hyperedges are formed as follows:
  • For the token Statistics, the LLM determines that $C_{1}$, $C_{4}$, and $C_{5}$ are semantically aligned in their scientific context, forming a hyperedge,
    $H_{1}=\{C_{1},C_{4},C_{5}\}\,.$
  • For the token Thermo,
    $\mathcal{G}(\text{Thermo})=\{C_{3},C_{4}\}\,,$
    and LLM groups these into another hyperedge,
    $H_{2}=\{C_{3},C_{4}\}.$
Using the hyperedges $H_{1}$ and $H_{2}$, we construct a hypergraph:
$H_{1}=\{C_{1},C_{4},C_{5}\},\quad H_{2}=\{C_{3},C_{4}\}.$
From this hypergraph, connected components are identified to group courses into standardized forms. For the given example, the connected component is
$\mathcal{C}_{1}=\{C_{1},C_{3},C_{4},C_{5}\}\,.$
Therefore, the raw courses $C_{1}$: Statistical Physics, $C_{3}$: Thermodynamics, $C_{4}$: Statistical Thermodynamics, and $C_{5}$: Statistical Mechanics are grouped into a single standardized course in the context of statistical physics and thermodynamics.
Appendix C Community detection in appropriate resolution
Figure 8: The PaI [26] values (left) and the average number of communities (right) as the functions of the resolution parameter $\gamma$ of the modularity function, as a result of $100$ independent runs of the Louvain algorithm [25] applied to the DSN in Sec. 3 for each $\gamma$ value. The points are the mean values, with the shades around them representing the standard deviation.
The resolution parameter $\gamma$ in the modularity function [37, 38] controls the overall scale of communities, where the modularity-maximization process with larger values of $\gamma$ tends to detect smaller communities. Although this flexibility provides a nice knob-like adjustable parameter for the scale of detected communities, it is nontrivial to decide which resolution yields the most meaningful or statistically reliable communities. In Ref. [26], the authors suggest a method to extract the range of $\gamma$ corresponding to the most self-consistent detection results, namely, the regions of locally minimum values of the partition inconsistency (PaI). We apply this method to our DSN with the Louvain algorithm [25], and as shown in Fig. 8, we find there values of $\gamma=0.56$, $1.12$, and $1.77$, corresponding to the locally consistent communities, so we take these values in addition to the trivial case $\gamma=0$ in Sec. 3 and Fig. 3.
Appendix D The log-log version of Fig. 5
To complement the linear-scale plots in Fig. 5, we also present the plots in the double-logarithmic scale in Fig. 9, to better highlight the tail behavior and sparsity of high-similarity discipline pairs.
Figure 9: The log-log version of Fig. 5. We exclude the zero values for visualization.
Appendix S1. The result of single and interdisciplinary curricula for ‘physics’ and ‘semiconductor engineering’
Table S1: We show the curriculum design process, beginning with the seed courses at the top and adding others with the criterion described in the main text (listed in the order of addition). The courses in boldface are common in all three curricula, highlighting their central role in integrating physics and semiconductor disciplines. The CSS value for this combination is 0.02017.
Physics Single Curriculum Physics & Semiconductor Interdisciplinary Curriculum Semiconductor Single Curriculum
Course Credits Course Credits Course Credits
Mathematical Physics 6 Mathematical Physics 6 Electromagnetism 6
Electromagnetism 6 Electromagnetism 6 Electronics 3
Modern Physics 3 Solid State Physics 3 Semiconductor Manufacturing 3
Solid State Physics 3 Modern Physics 3 Circuit Theory 6
Quantum Mechanics 6 Computational Physics 3 Semiconductor Devices 3
Computational Physics 3 Electronics 3 Microprocessor Fundamentals and Applications 3
Mechanics 6 Semiconductor Manufacturing 3 Applied Mathematics 3
Thermal Physics and
Statistical Mechanics
6 Circuit Theory 6 Semiconductor Physics 3
Optics 3 Semiconductor Devices 3 Semiconductor Engineering 3
Semiconductor Physics 3 Quantum Mechanics 6 Semiconductor Fabrication
and Laboratory Practice
3
Modern Physics and
Experimentation
3 Semiconductor Physics 3 Automation and Control Systems 3
Nuclear Physics 3 Mechanics 6 Engineering Mathematics 6
General Physics and
Experimentation
2 Optics 3 Digital Logic and Circuit Design 3
Advanced Physics 3 Thermal Physics and
Statistical Mechanics
6 Semiconductor Testing and Reliability 3
Nanophysics 3 Nuclear Physics 3 Advanced Semiconductor
Sensor Engineering and Applications
3
Biophysics 3 Modern Physics and Experimentation 3 Display Engineering 3
Statistical Physics 3 Microprocessor Fundamentals and Applications 3 Introduction to Semiconductor Engineering 3
Wave Optics and Physics 2 Engineering Mathematics 3 Semiconductor Packaging 3
Optics and Experiments 2 Semiconductor Engineering 3 Plasma Engineering 3
Particle Physics 3 Introduction to Semiconductor Engineering 3 Semiconductor Manufacturing Processes and Equipment 3
Electronics and Physics 3 Semiconductor Testing and Reliability 3 Digital Logic Design 3
Physics Education 3 Advanced Physics 3 Analog Integrated Circuit Design 3
Automation and Control Systems 3 Quantum Mechanics 3
Artificial Intelligence 3
Digital Circuit Design
and Semiconductor Technology
3
Chemistry of Semiconductors 3
Appendix S2. The result of single and interdisciplinary curricula for ‘chemistry’ and ‘high polymer engineering’
Table S2: We show the curriculum design process, beginning with the seed courses at the top and adding others with the criterion described in the main text (listed in the order of addition). The courses in boldface are common in all three curricula, highlighting their central role in integrating chemistry and polymer engineering disciplines. The CSS value for this combination is $0.03499$.
Chemistry Single Curriculum Chemistry & High Polymer Engineering Interdisciplinary Curriculum High Polymer Engineering Single Curriculum
Course Credits Course Credits Course Credits
Organic Chemistry 6 Polymer Chemistry 6 Polymer Chemistry 6
Inorganic Chemistry 6 Polymer Properties 3 Polymer Properties 3
Analytical Chemistry 6 Physical Chemistry 6 Physical Chemistry 6
Physical Chemistry 6 Organic Chemistry 6 Organic Chemistry 6
Instrumental Analysis 3 Inorganic Chemistry 6 Polymer Characterization and Analysis 3
Biochemistry 6 Analytical Chemistry 6 Functional Polymer Materials 3
Quantum Chemistry 3 Instrumental Analysis 3 Polymer Processing 3
Chemical Kinetics 3 Biochemistry 2 Polymer Science 3
Electrochemistry 3 Physical Chemistry
and Experimentation
2 Polymer Composites and Design 3
Organometallic Chemistry 3 Polymer Composites and Design 3 Polymer Nano Science and Engineering 3
Physical Organic Chemistry 3 Polymer Processing 3 Chemical Engineering Stoichiometry and Process Design 3
General Chemistry
and Laboratory
2 Polymer Physics and Chemistry 3 Polymer Physics and Chemistry 3
Nanochemistry 3 Polymer Rheology 3 Polymer Rheology 3
Physical Chemistry and Experimentation 2 Functional Polymer Materials 3 Analytical Chemistry 3
Biochemistry Experiments 2 Polymer Materials for Electronics and Information 3 Applied Mathematics 3
Polymer Science and Engineering 2 Instrumental Analysis 3
Biochemistry Experiments 3 Polymer Materials for Electronics and Information 3
Polymer Science 3 Polymer Science and Engineering 2
Unit Operations 3 Unit Operations 3
Engineering Mathematics 3 Engineering Mathematics 3
Electrochemistry 3 Polymer Properties Experimentation 2
Polymer Nano Science and Engineering 3 Biochemistry 2
Inorganic Chemistry 3
Polymer Surface and Interface Science 3
Polymer Physics 3
Appendix S3. The result of single and interdisciplinary curricula for ‘physics’ and ‘chemistry’
Table S3: We show the curriculum design process, beginning with the seed courses at the top and adding others with the criterion described in the main text (listed in the order of addition). The CSS value for this combination is 00.
Physics Single Curriculum Physics & Chemistry Interdisciplinary Curriculum Chemistry Single Curriculum
Course Credits Course Credits Course Credits
Mathematical Physics 6 Mathematical Physics 6 Organic Chemistry 6
Electromagnetism 6 Electromagnetism 6 Inorganic Chemistry 6
Modern Physics 3 Solid State Physics 3 Analytical Chemistry 6
Solid State Physics 3 Modern Physics 3 Physical Chemistry 6
Quantum Mechanics 6 Computational Physics 3 Inorganic Materials Science and Engineering 3
Computational Physics 3 Organic Chemistry 6 Instrumental Analysis 3
Mechanics 6 Inorganic Chemistry 6 Biochemistry 6
Thermal Physics and Statistical Mechanics 6 Physical Chemistry 6 Physical Chemistry and Experimentation 2
Optics 3 Analytical Chemistry 6 Polymer Chemistry 3
Semiconductor Physics 3 Inorganic Materials Science and Engineering 3 Quantum Chemistry 3
Modern Physics and Experimentation 3 Instrumental Analysis 6 Biochemistry Experiments 2
Nuclear Physics 3 Biochemistry 6 Nanochemistry 3
General Physics and Experimentation 2 Polymer Chemistry 3 General Chemistry and Laboratory 2
Advanced Physics 3 Quantum Mechanics 6 Chemical Kinetics 3
Nanophysics 3 Physical Chemistry and Experimentation 2 Electrochemistry 3
Biophysics 3 Mechanics 6 Organometallic Chemistry 3
Statistical Physics 3 Thermal Physics and Statistical Mechanics 6 Physical Organic Chemistry 3
Wave Optics and Physics 3 Optics 3
Optics and Experiments 2
Particle Physics 3
Electronics and Physics 3
Physics Education 3